# Large Language Models for Information Retrieval: A Survey

### v3

@@ 4.2.2 Generative Retriever
Traditional IR systems typically follow the “index-retrieval-rank” paradigm to locate relevant documents based on user queries, which has proven effective in practice. However, these systems usually consist of three separate modules: the index module, the retrieval module, and the reranking module. Therefore, optimizing these modules collectively can be challenging, potentially resulting in sub-optimal retrieval outcomes. Additionally, this paradigm demands additional space for storing pre-built indexes, further burdening storage resources. Recently, model-based generative retrieval methods [135, 136, 137] have emerged to address these challenges. These methods move away from the traditional “index-retrieval-rank” paradigm and instead use a unified model to directly generate document identifiers (i.e., DocIDs) relevant to the queries. In these model-based generative retrieval methods, the knowledge of the document corpus is stored in the model parameters, eliminating the need for additional storage space for the index. Existing methods have explored generating document identifiers through fine-tuning and prompting of LLMs [138, 139]
Fine-tuning LLMs. Given the vast amount of world knowledge contained in LLMs, it is intuitive to leverage them for building model-based generative retrievers. DSI [138] is a typical method that fine-tunes the pre-trained T5 models on retrieval datasets. The approach involves encoding queries and decoding document identifiers directly to perform retrieval. They explore multiple techniques for generating document identifiers and find that constructing semantically structured identifiers yields optimal results. In this strategy, DSI applies hierarchical clustering to group documents according to their semantic embeddings and assigns a semantic DocID to each document based on its hierarchical group. To ensure the output DocIDs are valid and do represent actual documents in the corpus, DSI constructs a trie using all DocIDs and utilizes a constraint beam search during the decoding process. Furthermore, this approach observes that the scaling law, which suggests that larger LMs lead to improved performance, is also applied to generative retrievers.
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. Based on the usage of LLMs, the existing LLM-based reranking methods can be divided into three paradigms: utilizing LLMs as supervised rerankers, utilizing LLMs as unsupervised rerankers, and utilizing LLMs for training data augmentation. These paradigms are summarized in Table V and will be elaborated upon in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [111]).
@@
Enc-only [140]
Enc-dec [13], [141], [142], [143]
Dec-only [131], [144], [145]
@@
Pointwise [146], [147], [148], [149], [150], [151]
Listwise [152], [153], [154]
Pairwise [155], [156]
Data Augmentation - [157], [158], [159], [160], [161], [162]
@@ 5.1.1 Encoder-only
The encoder-based rerankers represent a significant turning point in applying LLMs to document ranking tasks. They demonstrate how some pre-trained language models (e.g., BERT [59]) can be finetuned to deliver highly accurate relevance predictions. A representative approach is monoBERT [140], which transforms a query-document pair into a sequence “[CLS] query [SEP] document [SEP]” as the model input and calculates the relevance score by feeding the “[CLS]” representation into a linear layer. The reranking model is optimized based on the cross-entropy loss.
@@ 5.1.2 Encoder-Decoder
Although these generative loss-based methods outperform several strong ranking baselines, they are not optimal for reranking tasks. This stems from two primary reasons. First, it is commonly expected that a reranking model will yield a numerical relevance score for each query-document pair rather than text tokens. Second, compared to generation losses, it is more reasonable to optimize the reranking model using ranking losses (e.g., RankNet [163]). Recently, RankT5 [143] has directly calculated the relevance score for a query-document pair and optimized the ranking performance with “pairwise” or “listwise” ranking losses. An avenue for potential performance enhancement lies in the substitution of the base-sized T5 model with its larger-scale counterpart.
@@ 5.1.3 Decoder-only
Recently, there have been some attempts [131, 144, 145] to rerank documents by fine-tuning decoder-only models (such as LLaMA). For example, RankLLaMA [131] proposes formatting the query-document pair into a prompt “query: {query} document: {document} [EOS]” and utilizes the last token representation for relevance calculation. Besides, RankingGPT [144] has been proposed to bridge the gap between LLMs’ conventional training objectives and the specific needs of document ranking through two-stage training. The first stage involves continuously pretraining LLMs using a large number of relevant text pairs collected from web resources, helping the LLMs to naturally generate queries relevant to the input document. The second stage focuses on improving the model’s text ranking performance using high-quality supervised data and well-designed loss functions. Different from these pointwise rerankers [131, 144], Rank-without-GPT [145] proposes to train a listwise reranker that directly outputs a reranked document list. The authors first demonstrate that existing pointwise datasets (such as MS MARCO [111]), which only contain binary query-document labels, are insufficient for training efficient listwise rerankers. Then, they propose to use the ranking results of existing ranking systems (such as Cohere rerank API) as gold rankings to train a listwise reranker based on Code-LLaMA-Instruct.
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a query and a single document, and can be categorized into two types: relevance generation [146, 147] and query generation [148, 149, 150].
@@
where $S_{Y}$ and $S_{N}$ represent the LLM’s log-likelihood scores of “Yes” and “No” respectively. In addition to binary labels, Zhuang et al. [147] propose to incorporate fine-grained relevance labels (e.g., “highly relevant”, “somewhat relevant” and “not relevant”) into the prompt, which helps LLMs more effectively differentiate among documents with varying levels of relevance to a query.
@@
Although effective, these methods primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. As prompt is a key factor in instructing LLMs to perform various NLP tasks, it is important to optimize prompt for better performance. Along this line, a discrete prompt optimization method Co-Prompt [150] is proposed for better prompt generation in reranking tasks. Besides, PaRaDe [151] proposes a difficulty-based method to select few-show demonstrations to include in the prompt, proving significant improvements compared with zero-shot prompts.
Note that these pointwise methods rely on accessing the output logits of LLMs to calculate the query-document relevance scores. As a result, they are not applicable to closed-sourced LLMs, whose API-returned results do not include logits.
@@ 5.2.2 Listwise Methods
Although listwise methods have yielded promising performance, they still suffer from some weaknesses. First, according to the experimental results [152], only the GPT-4-based method can achieve competitive performance. When using smaller parameterized language models (e.g., FLAN-UL2 with 20B parameters), listwise methods may produce very few usable results and underperform many supervised methods. Second, the performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25 [152], revealing positional bias issues in the listwise ranking of LLMs. To alleviate this issue, Tang et al. [154] introduce a permutation self-consistency method, which involves shuffling the list in the prompt and aggregating the generated results to achieve a more accurate and unbiased ranking.
@@ 5.2.3 Pairwise Methods
In pairwise methods [155], LLMs are given a prompt that consists of a query and a document pair (see Figure 6 (c)). Then, they are instructed to generate the identifier of the document with higher relevance. To rerank all candidate documents, aggregation methods like AllPairs are used. AllPairs first generates all possible document pairs and aggregates a final relevance score for each document. To speed up the ranking process, efficient sorting algorithms, such as heap sort and bubble sort, are usually employed [155]. These sorting algorithms utilize efficient data structures to compare document pairs selectively and elevate the most relevant documents to the top of the ranking list, which is particularly useful in top-$k$ ranking. Experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than those typically employed in listwise methods (e.g., GPT3.5).
@@ 5.2.4 Comparison and Discussion
In this part, we will compare different unsupervised methods from various aspects to better illustrate the strengths and weaknesses of each method, which is summarized in Table VI. We choose representative methods [148, 146, 152, 155] in pointwise, listwise and pairwise ranking, and include several supervised methods [140, 13, 143] mentioned in Section 5.1 for performance comparison.
@@ 5.3 Utilizing LLMs for Training Data Augmentation
Furthermore, in the realm of reranking, researchers have explored the integration of LLMs for training data augmentation [157, 158, 159, 160, 161, 162]. For example, ExaRanker [157] generates explanations for retrieval datasets using GPT-3.5, and subsequently trains a seq2seq ranking model to generate relevance labels along with corresponding explanations for given query-document pairs. InPars-Light [158] is proposed as a cost-effective method to synthesize queries for documents by prompting LLMs. Contrary to InPars-Light [158], a new dataset ChatGPT-RetrievalQA [159] is constructed by generating synthetic documents based on LLMs in response to user queries.
@@ 5.4 Limitations
Although recent research on utilizing LLMs for document reranking has made significant progress, it still faces some challenges. For example, considering the cost and efficiency, minimizing the number of calls to LLM APIs is a problem worth studying. Besides, while existing studies mainly focus on applying LLMs to open-domain datasets (such as MSMARCO [111]) or relevance-based text ranking tasks, their adaptability to in-domain datasets [128] and non-standard ranking datasets [165] remains an area that demands more comprehensive exploration.
@@ 6 Reader
A naive strategy for implementing this function is to heuristically provide LLMs with documents relevant to the user queries or the previously generated texts to support the following generation. However, this passive approach limits LLMs to merely collecting documents from IR systems without active engagement. An alternative solution is to train LLMs to interact proactively with search engines. For example, LLMs can formulate their own queries instead of relying solely on user queries or generated texts for references. According to the way LLMs utilize IR systems in the reader module, we can categorize them into passive readers and active readers. Each approach has its advantages and challenges for implementing LLM-powered answer generation in IR systems. Furthermore, since the documents provided by upstream IR systems are sometimes too long to directly feed as input for LLMs, some compression modules are proposed to extractively or abstractively compress the retrieved contexts for LLMs to understand and generate answers for queries. We will present these reader and compressor modules in the following parts and briefly introduce the existing analysis work on retrieval-augmented generation strategy and their applications.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [178, 166, 168, 169, 170, 171, 179, 180, 175, 167, 173, 23, 176]. By this means, these approaches use the LLMs and IR systems separately, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 6.1.1 Once-Retrieval Reader
To obtain useful references for LLMs to generate responses for user queries, an intuitive way is to retrieve the top documents based on the queries themselves in the beginning. For example, REALM [166] adopts this strategy by directly attending the document contents to the original queries to predict the final answers based on masked language modeling. RAG [167] follows this strategy but applies the generative language modeling paradigm. However, these two approaches only use language models with limited parameters, such as BERT and BART. Recent approaches such as REPLUG [168] and Atlas [169] have improved them by leveraging LLMs such as GPTs, T5s, and LLaMAs for response generation. To yield better answer generation performances, these models usually fine-tune LLMs on QA tasks. However, due to the limited computing resources, many methods [170, 171, 179] choose to prompt LLMs for generation as they could use larger LMs in this way. Furthermore, to improve the quality of the generated answers, several approaches [181, 172] also try to train or prompt the LLMs to generate contexts such as citations or notes in addition to the answers to force LLMs to understand and assess the relevance of retrieved passages to the user queries. Some approaches [180] evaluate the importance of each retrieved reference using policy gradients to indicate which reference is more useful for generating. Besides, researchers explore instruction tuning LLMs such LLaMAs to improve their abilities to generate conclusive passages relying on retrieved knowledge [182, 183].

### v4

@@ 4.2.2 Generative Retriever
Traditional IR systems typically follow the “index-retrieval-rank” paradigm to locate relevant documents based on user queries. Though, this approach has proven its effective in practice, it usually consist of three separate modules: the index module, the retrieval module, and the reranking module. Optimizing these modules collectively can be challenging, potentially resulting in sub-optimal retrieval outcomes. Additionally, this paradigm demands additional storage space for pre-built indexes, further burdening storage resources. Recently, generative model-based retrieval methods [133, 134, 135] have emerged to address these challenges. These methods move away from the traditional “index-retrieval-rank” paradigm and instead use a unified model to directly generate document identifiers (i.e., DocIDs) relevant to the queries. In these generative retrieval methods, the knowledge of the document corpus is stored in the model parameters, eliminating the need for additional storage space for a separate index. Existing works have investigated the approaches of generating document identifiers through fine-tuning and prompting of LLMs [136, 137]
Fine-tuning LLMs. Given the vast amount of world knowledge contained in LLMs, it is intuitive to leverage them to build generative retrievers. DSI [136] is a typical method that fine-tunes the pre-trained T5 models on retrieval datasets. The approach involves encoding queries and decoding document identifiers directly to perform retrieval. They explore multiple techniques for generating document identifiers and find that constructing semantically structured identifiers yields optimal results. In this strategy, DSI applies hierarchical clustering to group documents according to their semantic embeddings and assigns a semantic DocID to each document based on its hierarchical group. To ensure the output DocIDs are valid and do represent actual documents in the corpus, DSI constructs a trie using all DocIDs and utilizes a constraint beam search during the decoding process. Furthermore, this approach observes that the scaling law, which suggests that larger LMs lead to improved performance, is also applied to generative retrievers. Though various generative retrievers have been proposed [136, 138, 139], most of them mainly focus on fine-tuning size-limited LMs on small-size document corpus (usually a subset of MSMARCO [99]). To analyze how the model size and document-corpus size impact the effectiveness of generative retrievers, Pradeep et al. [140] conducted a comprehensive analysis by scaling up corpus size from 100k to 8.8M and scaling model size up to 11B (T5-XXL). The primary findings are three-fold: (1) It is still challenging for generative retrievers to cover large-scale document corpus. (2) More model parameters often bring better performance. (3) Introducing synthetic queries generated from documents to expand training samples could significantly enhance the retrieval performance.
CorpusLM [141] further explores combining the retrieval and answering tasks together based on LLMs, making the two mutually reinforcing. Researchers devise various training tasks, e.g., DocID list generation, closed-book answers generation, and RAG generation, to sufficiently leverage and enhance the world knowledge of LLMs, improving the performance of these generation tasks.
@@
CorpusLM [141]
T5-base &
Llama2-7B-Chat
Generative Fine-tuning
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. According to the usage of LLMs, existing LLM-based reranking methods can be divided into three paradigms: utilizing LLMs as supervised rerankers, utilizing LLMs as unsupervised rerankers, and utilizing LLMs for training data augmentation. These paradigms are summarized in Table V and will be introduced in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [99]).
@@
Enc-only monoBERT [12]
Enc-dec
monoT5[13], Ju et al. [142], DuoT5 [143],
RankT5 [144], ListT5 [145]
Dec-only
RankLLaMA [122], TSARankLLM [146],
Q-PEFT [147],  Zhang et al. [148],
PE-Rank [149]
@@
Pointwise
Liang et al. [150], Zhuang et al. [151],
Guo et al. [152], Sachan et al. [153],
Zhuang et al. [154], Sun et al. [155],
Co-Prompt [156], DemoRank [157],
PARADE [158]
Listwise
RankGPT [159],  Ma et al. [160],
Tang et al. [161], TourRank [162],
Parry et al. [163]
Pairwise
PRP [164], Zhuang et al. [165],
PRP-Graph [166], Yan et al. [167]
Data
Augmentation
ExaRanker [168], ExaRanker-Open [169],
InPars-Light [170], Askari et al. [171],
Askari et al. [172], RankVicuna [173],
RankZephyr [174], Sun et al. [175]
@@ 5.1.1 Encoder-only
The encoder-based rerankers represent a significant turning point in applying LLMs to document ranking tasks. They demonstrate how some pre-trained language models (e.g., BERT [55]) can be fine-tuned to deliver highly accurate relevance predictions. A representative approach is monoBERT [12], which transforms a query-document pair into a sequence “[CLS] query [SEP] document [SEP]” as the model input and calculates the relevance score by feeding the “[CLS]” representation into a linear layer. The reranking model is optimized based on the cross-entropy loss.
@@ 5.1.2 Encoder-Decoder
Beyond the aforementioned studies, several studies have also explored different training losses and model architectures for reranker training. For example, RankT5 [144] is proposed to directly yield a numerical relevance score for each query-document pair during training and optimize the ranking performance with “pairwise” or “listwise” ranking losses. It differs significantly from the previous studies that optimize rerankers by generating text tokens and using a generation loss, and the use of ranking loss (e.g., RankNet [176]) is more reasonable and aligns better with the inherent nature of the ranking task. Besides,  Yoon et al. [145] propose ListT5, a listwise reranker based on Fusion-in-decoder architecture. It jointly takes multiple documents as input and directly generates a reranked document list during training and inference.
@@ 5.1.3 Decoder-only
Recently, there have been some attempts [122, 146, 147, 148, 149] to rerank documents by fine-tuning decoder-only models (such as LLaMA). For example, RankLLaMA [122] proposes formatting the query-document pair into a prompt “query: {query} document: {document} [EOS]” and utilizes the last token representation for relevance calculation. Besides, TSARankLLM [146] has been proposed to bridge the gap between LLMs’ conventional training objectives and the specific needs of document ranking through two-stage training. The first stage involves continuously pretraining LLMs using a large number of relevant text pairs collected from web resources, helping the LLMs to naturally generate queries relevant to the input document. The second stage focuses on improving the model’s text ranking performance using high-quality supervised data and well-designed loss functions.  Peng et al. [147] propose a query-dependent parameter efficient fine-tuning (Q-PEFT) approach for ranking, which helps the LLM generate true queries based on given documents. Different from these pointwise rerankers [122, 146, 147],  Zhang et al. [148] and  Liu et al. [149] proposes to train a listwise reranker that directly outputs a reranked document list. Specifically,  Zhang et al. [148] first demonstrate that existing pointwise datasets (such as MS MARCO [99]), which only contain binary query-document labels, are insufficient for training efficient listwise rerankers. Then, they propose to use the ranking results of existing ranking systems (such as Cohere rerank API) as gold rankings to train a listwise reranker based on Code-LLaMA-Instruct.  Liu et al. [149] propose PE-Rank, which compresses each document in the list into a single embedding and then inputs these document embeddings into reranker, which significantly reduces the input length and improves the efficiency of reranker.
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a query and a single document, and can be categorized into two types: relevance generation [157, 150, 151, 152] and query generation [153, 154, 156].
@@
where $S_{Y}$ and $S_{N}$ represent the LLM’s log-likelihood scores of “Yes” and “No” respectively. In addition to binary labels, Zhuang et al. [151] propose to incorporate fine-grained relevance labels (e.g., “highly relevant”, “somewhat relevant” and “not relevant”) into the prompt, which helps LLMs more effectively differentiate among documents with varying levels of relevance to a query.  Guo et al. [152] discuss the issues of inconsistent and biased relevance assessments of existing pointwise rerankers and introduce MCRanker that generates relevance scores based on a series of criteria from multiple perspectives.
@@
Although effective, these methods primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. Previous study [155] has shown that prompt has a significant impact on the performance of LLM reranker. Thus, how to design appropriate prompts for ranking task is an important problem. Along this line, a discrete prompt optimization method Co-Prompt [156] is proposed for better prompt generation in reranking tasks. Besides, PaRaDe [158] proposes a difficulty-based method to select the most difficult $k$ in-context demonstrations to include in the prompt, proving improvements compared with zero-shot performance. Nevertheless, the experiments in the paper indicate that such difficulty-based selection does not even show a significant advantage compared to random selection, showing that the demonstration selection in ranking task is a very challenging problem. The main challenge lies in the complex nature of query-document relationship, which requires effectively combining multiple demonstrations to help the LLM understand such relationship. Aiming to select more effective demonstrations for ranking task,  Liu et al. [157] propose DemoRank, an effective demonstration selection framework. The core component of DemoRank is a dependency-aware demonstration reranker, which reranks a list of demonstrations (usually obtained by a demonstration retriever) so that the combination of top-ranked demonstrations can yield better performance. An efficient method is proposed to construct the training samples for such demonstration reranker and a novel list-pairwise training loss is designed for optimization.
@@ 5.2.2 Listwise Methods
Although listwise methods have yielded promising performance, they still suffer from some weaknesses: (1) The performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25 [159], revealing positional bias issues in the listwise ranking of LLMs. (2) The use of sliding windows limits the number of documents that can be ranked each time, and the dependency between adjacent windows prevents parallelization of LLM inference, thereby reducing the efficiency of reranking. Recently, some studies have attempted to mitigate these issues. Tang et al. [161] introduce a permutation self-consistency method, which involves shuffling the list in the prompt and aggregating the generated results to achieve a more accurate and a positionally unbiased ranking.  Chen et al. [162] introduce a tournament mechanism into listwise ranking and propose TourRank, which parallelizes the reranking process through intelligent grouping and use a tournament-like points system to reduce the impact of the initial document order.  Parry et al. [163] propose a parallelizable partitioning algorithm for listwise ranking, which also aims at mitigating efficiency issues.  Reddy et al. [178] propose a novel listwise reranking approach which leverages the output logits of the first generated identifier to derive a ranked passage list.
@@ 5.2.3 Pairwise Methods
In pairwise methods, LLMs are given a prompt that consists of a query and a document pair (see Figure 5 (c)). Then, they are instructed to generate the identifier of the document with higher relevance. To rerank all candidate documents, aggregation methods like AllPairs are used [164]. AllPairs first generates all possible document pairs, yields discrete judgments for each pair (e.g., Document 1 or Document 2), and aggregates a final relevance score for each document. Efficient sorting algorithms, such as heap sort and bubble sort, are employed to speed up the ranking process. These sorting algorithms utilize efficient data structures to compare document pairs selectively and elevate the most relevant documents to the top of the ranking list, which is particularly useful in top-$k$ ranking. Experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than those typically employed in listwise methods (e.g., GPT3.5). Building on the pairwise prompting approach,  Luo et al. [166] introduce an innovative scoring unit that leverages the generation probability of judgments instead of discrete judgments, and further design a graph-based aggregation approach to obtain a final relevance score for each document. Besides, the pairwise comparison can also be utilized as a post-processing step to adjust the relevance scores generated by the pointwise LLM reranker [167].
@@ 5.2.4 Comparison and Discussion
We compare different unsupervised methods from various aspects to better illustrate the strengths and weaknesses of each method, which is summarized in Table VI. The representative methods [153, 150, 159, 164] in pointwise, listwise, and pairwise ranking, and some supervised methods [12, 13, 144] mentioned in Section 5.1 are selected for performance comparison.
@@ 5.3 Utilizing LLMs for Training Data Augmentation
Furthermore, in the realm of reranking, researchers have explored the integration of LLMs for training data augmentation [168, 170, 171, 173, 174, 175]. For example, ExaRanker [168] and ExaRanker-Open [169] generate explanations for query-passage pairs using GPT-3.5 and open-source LLMs respectively, and subsequently trains a seq2seq ranking model to generate relevance labels along with corresponding explanations. InPars-Light [170] is proposed as a cost-effective method to synthesize queries for documents by prompting LLMs.  Askari et al. [171] proposes to generate synthetic documents based on LLMs in response to user queries. Furthermore,  Askari et al. [172] propose to utilize reinforcement learning to improve the quality of synthetic documents generated by LLMs.
@@ 5.4 Limitations
Although recent research on utilizing LLMs for document reranking has made significant progress, it still has some limitations. First, due to the reliance on API calls and a large number of parameters, the process of LLM ranking could be expensive and inefficient. Therefore, achieving a trade-off between the cost/efficiency and performance of LLMs is a topic worth discussing. Along this line,  Rashid et al. [179] propose a budget-aware ranking solution which maximizes the LLM’s performance within a given budget. Besides,  Meng et al. [180] systematically discuss the improvements in ranking efficiency and effectiveness brought by the rank list truncation technique. Second, while existing studies mainly focus on applying LLMs to open-domain datasets (such as MSMARCO [99]) or relevance-based text ranking tasks, their adaptability to in-domain datasets [113] and non-standard ranking datasets [181] remains an area that demands more comprehensive exploration.
@@ 6 Reader
A naive strategy for implementing this function is to heuristically provide LLMs with documents relevant to the user queries or the previously generated texts to support the following generation. However, this passive approach limits LLMs to merely collecting documents from IR systems without active engagement. An alternative solution is to train LLMs to interact proactively with search engines. For example, LLMs can formulate their own queries instead of relying solely on user queries or generated texts for references. According to the way LLMs utilize IR systems in the reader module, we can categorize them into passive readers and active readers. Each approach has its advantages and challenges for implementing LLM-powered answer generation in IR systems. Furthermore, since the documents provided by upstream IR systems are sometimes too long to directly feed as input for LLMs, some compression modules are proposed to extractively or abstractively compress the retrieved contexts for LLMs to understand and generate answers for queries. We will present these reader and compressor modules in the following parts and briefly introduce the existing analysis work on retrieval-augmented generation (RAG) strategy and their applications.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [194, 182, 184, 185, 186, 187, 195, 196, 191, 183, 189, 23, 192, 197]. By this means, these approaches use the LLMs and IR systems separately, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 6.1.1 Once-Retrieval Reader
To obtain useful references for LLMs to generate responses for user queries, an intuitive way is to retrieve the top documents based on the queries themselves in the beginning. For example, REALM [182] adopts this strategy by directly attending the document contents to the original queries to predict the final answers based on masked language modeling. RAG [183] follows this strategy but applies the generative language modeling paradigm. However, these two approaches only use language models with limited parameters, such as BERT and BART. Recent approaches such as REPLUG [184] and Atlas [185] have improved them by leveraging LLMs such as GPTs, T5s, and LLaMAs for response generation. To yield better answer generation performances, these models usually fine-tune LLMs on QA tasks. However, due to the limited computing resources, many methods [186, 187, 195, 198] choose to prompt LLMs for generation as they could use larger LMs in this way. Furthermore, to improve the quality of the generated answers, several approaches [199, 188] also try to train or prompt the LLMs to generate contexts such as citations or notes in addition to the answers to force LLMs to understand and assess the relevance of retrieved passages to the user queries. ActiveRAG [200] and PG-RAG [201] improve them by using knowledge construction during the answer generation process. Some approaches [196, 202] evaluate the importance of each retrieved reference using policy gradients to indicate which reference is more useful for generating. Specifically,  [202] utilize LLMs themselves to provide importance for different references which also supply additional training signals. Besides, researchers explore instruction tuning LLMs such LLaMAs to improve their abilities to generate conclusive passages relying on retrieved knowledge [203, 204, 205]. During the training of LLM-based readers, some approaches [206] explore the strategy of contrastive learning by augmenting training data by removing and replacing retrieved passages to improve the generating performances. Additionally, SPRING [207] inserts several trainable tokens between the retrieved documents and issued questions for better optimization of the reader. R2AG [208] extracts features from retrieval models and attaches them to the reference contents to overcome the semantic gaps between LLMs and retrievers.  Yoran et al. [209] also propose to generate noisy training data to help LLMs generate correct answers while irrelevant contents are included in the retrieved contexts. RAAT [210] and ATM [211] further solve the noisy problem by introducing the adversarial training strategy.